# Internet of Agents: Fundamentals, Applications, and Challenges

### v1

@@ I Introduction
The rapid advancement of large models, including large language models (LLMs) and vision-language models (VLMs), has ushered in a new era of artificial intelligence (AI) agents (or agentic AI) [1, 2], transforming them from isolated, task-specific models into autonomous, interactive entities. These agents can perceive, reason, and act independently, capable of seamless collaboration with humans and other agents in complex environments, marking a pivotal step toward artificial general intelligence (AGI) [3, 4]. From virtual assistants to physically embodied systems such as humanoid robots, autonomous unmanned aerial vehicles (UAVs), and intelligent vehicles, AI agents are rapidly weaving into daily life [5]. Tech giants have declared their adventure to develop next-generation AI agents, such as OpenAI Operator and ByteDance’s Doubao Agent. For instance, platforms such as Hugging Face host over 1 million open-source models, while Tencent Yuanbao support over 100K specialized agents. According to Gartner [6], by 2028, at least 15% of daily tasks will be autonomously performed by AI agents, and 33% of enterprise applications will incorporate agent-driven intelligence. As AI agents proliferate, they are poised to act as “new citizens” in digital and physical spaces, reshaping economic structures and human social interactions.
The widespread adoption of agents has spurred the need for real-time cross-domain agent communication and coordination, particularly in scenarios such as smart cities including millions of heterogeneous agents. The Internet of agents (IoA) [7, 8], also referred to as the agentic web, emerges as a foundational infrastructure for next-generation intelligent systems that enables seamless interconnection, autonomous agent discovery, dynamic task orchestration, and collaborative reasoning among large-scale virtual/embodied agents. Unlike the human-centric Internet, IoA is agent-centric and prioritizes inter-agent interactions [9], where the exchanged information shifts from human-oriented data (e.g., text, images, and audio) to machine-oriented data objects (e.g., model parameters, encrypted tokens, and latent representations). Furthermore, interaction methods are evolving beyond graphical user interfaces (GUIs) toward semantic-aware and goal-driven communications [10] via auto-negotiation [11]. Additionally, by offering scalable networked AI inference and shared sensing capabilities, IoA empowers resource-constrained agents, e.g., mobile devices and UAVs, with access to advanced AI capabilities [12] and beyond-line-of-sight (BLOS) perception. By fostering seamless interoperability between virtual and embodied agents from distinct entities, IoA introduces new connection paradigms and traffic patterns beyond the current Internet and mobile Internet. As this ecosystem continues to evolve, IoA is poised to become the backbone of a new era of human-AI symbiosis, reshaping global connectivity and pushing the frontiers of intelligent systems.
@@
  • Agent-Native Interface: Current computer-use agents (e.g., OpenAI’s Operator) rely on mimicking human GUI actions (e.g., clicks and keystrokes) to control browsers and apps [14], incurring high screen-scraping overhead. IoA should empower agents to interact natively (e.g., APIs or semantic communication protocols) with other agents and Internet resources [9, 13], rather than emulating human behavior.
@@ I-A Comparison with Existing Surveys and Contributions of Our Work
Recent research in large model-based agents and multi-agent systems has attracted considerable attention across both academia and industry. Jin et al. [15] review intelligent decision-making approaches, algorithms, and models in MAS and categorize them into rule-based, game theoretical, evolutionary algorithm-based, multi-agent reinforcement learning (MARL)-based, and LLM-based methods. Guo et al. [16] systematically examine the development of LLM-based MAS by addressing the agent-environment interface, LLM agent characterization, inter-agent communication strategies, and capability acquisition paradigms, while also highlighting applications in problem-solving and world simulation. Tran et al. [17] classify LLM-based multi-agent collaborative systems based on key characteristics including type, strategy, structure, and coordination. Li et al. [18] offer a comprehensive survey on the construction of LLM-based MAS with a focus on problem-solving and world simulation. Wang et al. [2] review the overall framework of LLM agents, covering enabling technologies, core characteristics, collaboration paradigms, and their associated security and privacy challenges, along with potential defense strategies in future AI agent systems. Wu et al. [19] present a review of LLM-based multi-agent autonomous driving systems by addressing multivehicle interactions, vehicle-infrastructure communication, and human-vehicle codriving. He et al. [20] systematically assess the capabilities and limitations of LLM-based MAS applications in software engineering. Amirkhani et al. [21] provide an overview of consensus in MAS by reviewing taxonomies, dynamic models, protocols, control mechanisms, and applications.
Existing surveys mainly focus on multi-agent systems (MAS), which have three challenges. 1) Ecosystem isolation: Existing frameworks limit agents only to their own environments, restricting the integration of third-party agents and reducing the diversity of capabilities. 2) Single-device simulation: Most MAS are confined to single-device simulations, which contrast sharply with real-world scenarios where agents operate across multiple devices and varied geographic locations. 3) Rigid communication and coordination: Most of existing agent protocols and state transitions are hard-coded and fail to capture the dynamic, task-specific nature of practical collaboration. In contrast, this survey focuses on the networking aspects of large model-based agents, addressing their architectures and open challenges of the Internet of agents (IoA). Table II compares the contributions of our survey with previous related surveys in the field of IoA.
@@
We provide a comprehensive survey on the fundamentals, emerging applications, and open challenges of IoA, paving the way for its further development and practical deployment. Our survey aims to: 1) provide a comprehensive understanding of the IoA architecture and its networking mechanisms that facilitate the integration of diverse, distributed agents; and 2) critically examine the operational paradigms of IoA, including dynamic agent discovery, task orchestration, communication protocols, and consensus mechanisms. The main contributions include:
  • General IoA Architecture. We provide a comprehensive overview of the IoA architecture, discussing its hierarchical framework, distinguishing features from the traditional Internet, key characteristics, and emerging applications.
  • Key Enablers of IoA. We analyze the building blocks and enabling technologies underpinning IoA, including capability notification and discovery, dynamic task orchestration and matching, adaptive communication protocols, consensus and conflict resolution mechanisms, and economic models.
  • Open Challenges of IoA. We identify several unresolved issues in the IoA domain and point out future research directions to foster its broad adoption and robust development.
@@ II-A Architecture of IoA
1) Agent Types. In IoA, agents can be broadly categorized into two principal types: virtual agents and embodied agents.
@@
  • Memory. The memory module underpins an agent’s ability to learn, adapt, and personalize its behavior by maintaining both short-term context (e.g., ongoing dialogues and recent task results) and long-term knowledge (e.g., accumulated domain facts and past interaction records) [3]. By combining real-time updates with historical data, it enables continual learning, personalization, and informed decision-making. Memory can be organized into three types. (i) Short-term memory [29] is a limited context buffer for recent observations and dialogue turns, ensuring coherent, up-to-date reasoning. (ii) Long-term memory [30] is external vector stores or knowledge bases that archive historical experiences, domain facts, and policies, accessed via retrieval-augmented generation (RAG) [31] for informed planning. (iii) Hybrid memory [32] is a dynamic cache that promotes key short-term fragments into long-term storage and retrieves relevant long-term entries back into the working buffer.
@@
  • Action. The action module brings an agent’s plans to life by interfacing with both physical actuators and software tools. It also monitors execution outcomes via the perception module, adjusting behaviors to handle unexpected obstacles or changing objectives. It encompasses two main capabilities: (i) Embodied operations: translates high-level plans into real-world behaviors such as robotic grasping, navigation, or manipulation. For instance, in SayCan [35], a warehouse robot can use an LLM to suggest pickup poses (“Say”) and a learned affordance model to verify feasibility (“Can”), ensuring reliable execution under changing conditions. (ii) Tool invocation & creation: empowers agents to call external services such as search engines, APIs, or databases to gather information or execute subtasks. Agents can also synthesize new tools (e.g., custom data‐processing scripts) to expand their action repertoire for improved efficiency.
Figure 3: An overview of IoA encompassing heterogeneous autonomous agents across diverse domains, with an illustrative example of cross-domain agent interactions. Consider two agents operating in distinct domains, e.g., domains 1 and 4. Initially, each agent registers with a proxy agent hosted on its respective domain gateway to facilitate agent discovery. Gateway proxy agents then broadcast the capabilities of all registered agents within their domain. Upon identifying a common task, the two agents initiate collaboration, facilitated by relay gateways that manage inter-domain communication. Throughout the interaction, gateway proxy agents coordinate and synchronize the agents’ states, ensuring consistency and coherence across domains.
@@
3) Interconnected Sub-IoA. As illustrated in Fig. 3, the IoA architecture comprises multiple interconnected and domain-specific agent networks, referred to as sub-IoAs. Each sub-IoA operates semi-autonomously and is anchored by a designated gateway node that hosts a gateway proxy agent [8], responsible for orchestrating resources, agent egister & discovery, agent capacity notification within its domain, as well as cross-domain relays and state synchronization. This hierarchical structure enables scalable, modular deployment across diverse environments.
@@ II-B Key Characteristics of IoA
2) High Dynamics. IoA networks are inherently dynamic, characterized by the on-demand creation, migration, and termination of virtual agents, as well as the physical mobility of embodied agents such as UAVs and robots. This agility enables real-time reconfiguration of agent teams to accommodate varying task demands, environmental conditions, or resource constraints. Additionally, agent’s capabilities, as well as task orchestration and interaction protocols, adjust continuously.
@@
  • Capability Variance: Agents range from battery-constrained end devices to large model agents with abundant GPU and memory. Task assignments and workload offloading should dynamically adapt to agents’ resource constraints.
@@
  • Hierarchical Sharding: Agents self-organize into hierarchical clusters or “shards” based on geography, function, or task affinity [18]. For instance, domain gateways host proxy agents that register local agents, broadcast their capabilities, and coordinate cross-domain collaboration and state synchronization via relay gateways, as shown in Fig. 3.
@@
This shift from data-centric to meaning-aware communication is fundamental to building goal-oriented, adaptive, and scalable multi-agent systems.
TABLE III: Key Comparisons between Traditional Internet, IoT, and IoA
@@ II-C Key Differences Among Internet, IoT, and IoA
IoA represents a paradigm shift from the traditional Internet and IoT by fundamentally redefining core objectives, architectural principles, operational mechanisms, and interaction patterns, as summarized in Table III.
@@ II-D Emerging IoA Application Scenarios
1) Typical Multi-Agent Frameworks: Recent advancements in LLMs have spurred the development of various agent frameworks capable of autonomous or semi-autonomous task execution through advanced reasoning, planning, and tool integration. AutoGPT [52] implements autonomous goal-driven agents by iteratively refining actions through recursive self-prompting. AutoGen [41] introduces conversational agent networks where multiple LLM-powered agents (e.g., assistants and user proxies) interact dynamically, supporting both human-in-the-loop and fully automated workflows. LangChain [53] provides a modular framework for chaining LLM calls with external tools (e.g., APIs and databases) and memory systems, offering flexibility for customizable agent pipelines. MetaGPT [37] simulates software development teams with role-based agents (e.g., product managers and engineers) that follow structured workflows, enhancing multi-agent collaboration. Other notable frameworks include BabyAGI [54], which focuses on task-driven autonomous iteration, and CAMEL [47], which explores role-playing agent societies with communicative agents. Table IV shows the comparisons of mainstream open-source multi-agent frameworks.
TABLE IV: Comparison of Mainstream Agent Frameworks
@@
MetaGPT[37] Role-based agents, structured workflows, multi-agent collaboration High efficiency in complex tasks Require predefined roles, less flexible Automated software development
LangChain[53] Modular LLM chaining, memory, tool integration Highly customizable, supports diverse tools Steeper learning curve, manual tuning needed Custom agent pipelines, RAG systems
AutoGPT[52] Recursive self-prompting, goal-driven autonomy Fully autonomous task execution Prone to loops, high compute cost General task automation
AutoGen[41] Conversational MAS, human-in-the-loop support Flexible collaboration, dynamic interactions Complex setup for optimal performance Multi-agent dialogue, AI teamwork
BabyAGI[54] Task-driven autonomous iteration, simple architecture Lightweight, easy to deploy Limited reasoning depth Small-scale automation
CAMEL[47] Role-playing agent societies, communicative agents Simulates human-like interactions Less optimized for tool use Social AI, research simulations
@@
2) Roadmap: Existing representative industrial practices include computer use, AI smartphones, Anthropic’s model context protocol (MCP) [13], and Google’s agent-to-agent (A2A) Protocol [9]. (i) Computer-Use Agents (CUAs): As existing Internet services are originally built for humans, CUAs [14] such as OpenAI’s Operator mimic human GUI interactions (e.g., mouse clicks and keystrokes) to drive browsers and mobile interfaces, but incur significant screen-scraping overhead that undermines agent efficiency. (ii) AI Smartphones: AI smartphones (e.g., Apple Intelligence) enable cross-App data access and function calls by exposing App APIs to agents, better aligned with agent’s strength in processing raw data than GUI emulation. (iii) MCP places large models at the core of the ecosystem, treating the entire Internet as its contextual knowledge base and toolkit. Standardized APIs, dynamic plugin support, and persistent context management allow a single model to orchestrate across diverse services. (iv) A2A protocol aims to form a decentralized agent network where agents autonomously discover, negotiate, and collaborate with each other. Ultimately, IoA should provide (1) seamless interoperability to eliminate data silos, (2) agent-native interfaces for direct efficient inter-agent interactions, and (3) protocol-driven self-organizing cooperation among agents.
@@
3) Applications: The IoA paradigm holds transformative potential across diverse sectors. As shown in Fig. 5, we illustrate its applications in three representative scenarios: smart homes, smart factories, and smart cities.
@@
  • Scenario 2: Coordination between Agents in an IoA Subnet and and External Agents in Smart Factory. Within a smart factory, the IoA subnet connects on-site manufacturing agents (e.g., robotic arms, autonomous guided vehicles, and digital twin controllers) with external ecosystem participants (such as component suppliers, logistics UAVs, and cloud analytic agents). For instance, an online production supervisor agent can request spare parts through a supplier’s digital assistant, assign an autonomous forklift to negotiate pickup timing with a delivery UAV, and authenticate identities via QR-code or device-to-device (D2D) communications. Throughout this process, the network provides digital identity verification and sensing-based path-planning services, ensuring that materials arrive just-in-time and production workflows adapt responsively to supply-chain dynamics.

### v2

@@ I Introduction
The rapid advancement of large models, including large language models (LLMs) and vision-language models (VLMs), has ushered in a new era of artificial intelligence (AI) agents (or agentic AI) [1, 2], transforming them from isolated, task-specific models into autonomous, interactive entities. These agents can perceive, reason, and act independently, capable of seamless collaboration with humans and other agents in complex environments, marking a pivotal step toward artificial general intelligence (AGI) [3, 4]. From virtual assistants to physically embodied systems such as humanoid robots, autonomous unmanned aerial vehicles (UAVs), and intelligent vehicles, AI agents are rapidly weaving into daily life [5]. Tech giants have declared their adventure to develop next-generation AI agents, such as OpenAI Operator and ByteDance's Doubao Agent. For instance, platforms such as Hugging Face host over 1 million open-source models, while Tencent Yuanbao support over 100K specialized agents. According to Gartner [6], by 2028, at least 15% of daily tasks will be autonomously performed by AI agents, and 33% of enterprise applications will incorporate agent-driven intelligence. As AI agents proliferate, they are poised to act as ``new citizens" in digital and physical spaces, reshaping economic structures and human social interactions.
The widespread adoption of agents has spurred the need for real-time cross-domain agent communication and coordination, particularly in scenarios such as smart cities including millions of heterogeneous agents. The Internet of agents (IoA) [7, 8] emerges as a foundational infrastructure for next-generation intelligent systems that enables seamless interconnection, autonomous agent discovery, dynamic task orchestration, and collaborative reasoning among large-scale virtual/embodied agents. Unlike the human-centric Internet, IoA is agent-centric and prioritizes inter-agent interactions [9], where the exchanged information shifts from human-oriented data (e.g., text, images, and audio) to machine-oriented data objects (e.g., model parameters, encrypted tokens, and latent representations). Furthermore, interaction methods are evolving beyond graphical user interfaces (GUIs) toward semantic-aware and goal-driven communications [10] via auto-negotiation [11]. Additionally, by offering scalable networked AI inference and shared sensing capabilities, IoA empowers resource-constrained agents, e.g., mobile devices and UAVs, with access to advanced AI capabilities [12] and beyond-line-of-sight (BLOS) perception. By fostering seamless interoperability between virtual and embodied agents from distinct entities, IoA introduces new connection paradigms and traffic patterns beyond the current Internet and mobile Internet. As this ecosystem continues to evolve, IoA is poised to become the backbone of a new era of human-AI symbiosis, reshaping global connectivity and pushing the frontiers of intelligent systems.
@@
  • Agent-Native Interface: Current computer-use agents (e.g., OpenAI's Operator) rely on mimicking human GUI actions (e.g., clicks and keystrokes) to control browsers and apps [14], incurring high screen-scraping overhead. IoA should empower agents to interact natively (e.g., APIs or semantic communication protocols) with other agents and Internet resources [9, 13], rather than emulating human behavior.
@@ I-A Comparison with Existing Surveys and Contributions of Our Work
Recent research in large model-based agents and MAS has attracted considerable attention across both academia and industry. Jin et al. [15] review intelligent decision-making approaches, algorithms, and models in MAS and categorize them into rule-based, game theoretical, evolutionary algorithm-based, multi-agent reinforcement learning (MARL)-based, and LLM-based methods. Guo et al. [16] systematically examine the development of LLM-based MAS by addressing the agent-environment interface, LLM agent characterization, inter-agent communication strategies, and capability acquisition paradigms, while also highlighting applications in problem-solving and world simulation. Tran et al. [17] classify LLM-based multi-agent collaborative systems based on key characteristics including type, strategy, structure, and coordination. Li et al. [18] offer a comprehensive survey on the construction of LLM-based MAS with a focus on problem-solving and world simulation. Wang et al. [2] review the overall framework of LLM agents, covering enabling technologies, core characteristics, collaboration paradigms, and their associated security and privacy challenges, along with potential defense strategies in future AI agent systems. Wu et al. [19] present a review of LLM-based multi-agent autonomous driving systems by addressing multivehicle interactions, vehicle-infrastructure communication, and human-vehicle codriving. He et al. [20] systematically assess the capabilities and limitations of LLM-based MAS applications in software engineering. Amirkhani et al. [21] provide an overview of consensus in MAS by reviewing taxonomies, dynamic models, protocols, control mechanisms, and applications.
Existing surveys mainly focus on MAS, which have three challenges. 1) Ecosystem isolation: Existing frameworks limit agents only to their own environments, restricting the integration of third-party agents and reducing the diversity of capabilities. 2) Single-device simulation: Most MAS are confined to single-device simulations, which contrast sharply with real-world scenarios where agents operate across multiple devices and varied geographic locations. 3) Rigid communication and coordination: Most of existing agent protocols and state transitions are hard-coded and fail to capture the dynamic, task-specific nature of practical collaboration. In contrast, this survey focuses on the networking aspects of large model-based agents, addressing their architectures and open challenges of the Internet of agents (IoA). Table II compares the contributions of our survey with previous related surveys in the field of IoA.
@@
We provide a comprehensive survey on the fundamentals, emerging applications, and open challenges of IoA, paving the way for its future development. Our survey aims to: 1) provide a comprehensive understanding of the IoA architecture and its networking mechanisms that facilitate the integration of diverse, distributed agents; and 2) critically examine the operational paradigms of IoA, including dynamic agent discovery, task orchestration, communication protocols, and consensus mechanisms. The main contributions include:
  • General IoA Architecture. We provide a holistic overview of the IoA architecture, outlining its hierarchical framework, distinguishing features from the traditional Internet, key characteristics, and emerging applications.
  • Key Enablers of IoA. We analyze critical enablers that support IoA, including capability notification & discovery, dynamic task orchestration & matching, adaptive communication protocols, consensus & conflict resolution mechanisms, economic models, and regulation models.
  • Open Challenges of IoA. We identify unresolved issues in the IoA domain and point out future research directions to foster its broad adoption and sustainable evolution.
@@ II-A Architecture of IoA
1) Agent Types. Unlike traditional symbolic agents that rely on rule-based or logic-driven mechanisms for decision-making, LLM-powered agents employ generative models for reasoning, planning, and natural language interaction. In addition, robotic or embodied agents operate in the physical world by coupling perception with actuation. Table III summarizes the typology of agents. Overall, within the IoA, agents can be broadly classified into two principal categories: virtual agents and embodied agents.
@@
TABLE III: Comparison Among Symbolic Agents, LLM-powered Virtual Agents, and Robotic/Embodied Agents
Agent Type Core Mechanism Strengths Limitations Roles in IoA
Symbolic Agents Logic-based or rule-driven inference and decision-making High interpretability, formal guarantees, efficient in structured domains Limited scalability, less flexible in unstructured environments Protocol enforcement, rule-based coordination, compliance checking
LLM-powered Virtual Agents Generative large models for reasoning, planning, and natural language interaction Strong adaptability, contextual reasoning, multi-domain capability Computationally expensive, unpredictable behavior, reliance on data quality Conversational agents, code generation, research assistants
Robotic/Embodied Agents Coupling perception with actuation in the physical world Direct real-world interaction, physical task execution Hardware constraints, safety/reliability concerns, limited reasoning ability Manufacturing, autonomous vehicles, healthcare robotics
@@
  • Memory. The memory module underpins an agent's ability to learn, adapt, and personalize its behavior by maintaining both short-term context (e.g., ongoing dialogues and recent task results) and long-term knowledge (e.g., accumulated domain facts and past interaction records) [3]. By combining real-time updates with historical data, it enables continual learning, personalization, and informed decision-making. Memory can be organized into three types. (i) Short-term memory [29] is a limited context buffer for recent observations and dialogue turns, ensuring coherent, up-to-date reasoning. (ii) Long-term memory [30] is external vector stores or knowledge bases that archive historical experiences, domain facts, and policies, accessed via retrieval-augmented generation (RAG) [31] for informed planning. (iii) Hybrid memory [32] is a dynamic cache that promotes key short-term fragments into long-term storage and retrieves relevant long-term entries back into the working buffer.
@@
  • Action. The action module brings an agent's plans to life by interfacing with both physical actuators and software tools. It also monitors execution outcomes via the perception module, adjusting behaviors to handle unexpected obstacles or changing objectives. It encompasses two main capabilities: (i) Embodied operations: translates high-level plans into real-world behaviors such as robotic grasping, navigation, or manipulation. For instance, in SayCan [35], a warehouse robot can use an LLM to suggest pickup poses (``Say'') and a learned affordance model to verify feasibility (``Can''), ensuring reliable execution under changing conditions. (ii) Tool invocation & creation: empowers agents to call external services such as search engines, APIs, or databases to gather information or execute subtasks. Agents can also synthesize new tools (e.g., custom data‐processing scripts) to expand their action repertoire for improved efficiency.
Figure 3: An overview of IoA encompassing heterogeneous autonomous agents across diverse domains, with an illustrative example of cross-domain agent interactions. Consider two agents operating in distinct domains, e.g., domains 1 and 4. Initially, each agent registers with a proxy agent hosted on its respective domain gateway to facilitate agent discovery. Gateway proxy agents then broadcast the capabilities of all registered agents within their domain. Upon identifying a common task, the two agents initiate collaboration, facilitated by relay gateways that manage inter-domain communication. Throughout the interaction, gateway proxy agents coordinate and synchronize the agents' states, ensuring consistency and coherence across domains.
@@
3) Interconnected Sub-IoA. As illustrated in Fig. 3, the IoA architecture comprises multiple interconnected and domain-specific agent networks, referred to as sub-IoAs. Each sub-IoA operates semi-autonomously and is anchored by a designated gateway node that hosts a gateway proxy agent [8], responsible for orchestrating resources, agent register & discovery, agent capacity notification within its domain, as well as cross-domain relays and state synchronization. This hierarchical structure enables scalable, modular deployment across diverse environments.
@@ II-B Key Characteristics of IoA
2) High Dynamics. IoA networks are inherently dynamic, characterized by the on-demand creation, migration, and termination of virtual agents, as well as the physical mobility of embodied agents such as UAVs and robots. This agility enables real-time reconfiguration of agent teams to accommodate varying task demands, environmental conditions, or resource constraints. Additionally, agent's capabilities, as well as task orchestration and interaction protocols, adjust continuously.
@@
  • Capability Variance: Agents range from battery-constrained end devices to large model agents with abundant GPU and memory. Task assignments and workload offloading should dynamically adapt to agents' resource constraints.
@@
  • Hierarchical Sharding: Agents self-organize into hierarchical clusters or ``shards'' based on geography, function, or task affinity [18]. For instance, domain gateways host proxy agents that register local agents, broadcast their capabilities, and coordinate cross-domain collaboration and state synchronization via relay gateways, as shown in Fig. 3.
@@
This shift from data-centric to meaning-aware communication is fundamental to building goal-oriented, adaptive, and scalable MAS.
TABLE IV: Key Comparisons between Traditional Internet, IoT, and IoA
@@
TABLE V: Conceptual Comparison of IoA with Related Paradigms
Paradigm Scope & Focus Scale Autonomy & Adaptivity Typical Applications
MAS Coordinated problem-solving in bounded environments Moderate, domain-specific Agents collaborate under centralized/semi-centralized orchestration Home robots coordination, software development
Agentic Web Intelligent services on top of the Web Large, web-centric Enhancing user services via intelligent agents Personalized assistants, intelligent web navigation
Web 3.0 Decentralized, semantic, and ownership-driven web Internet-wide Limited autonomy; emphasis on data interoperability Decentralized finance, semantic knowledge sharing
IoA Internet-like fabric interconnecting heterogeneous agents Large-scale, cross-domain High autonomy, dynamic adaptivity, open collaboration Cross-domain agent ecosystems
@@ II-C Key Differences Among Internet, IoT, and IoA
IoA represents a paradigm shift from the traditional Internet and IoT by fundamentally redefining core objectives, architectural principles, operational mechanisms, and interaction patterns, as summarized in Table IV.
@@
While the IoA is inspired by earlier paradigms, it is distinct in scope and objectives. Traditional MAS primarily focuses on coordinated problem-solving within bounded environments, often under centralized or semi-centralized orchestration. In contrast, the IoA emphasizes large-scale, open, and heterogeneous ecosystems where autonomous agents interact dynamically across domains, infrastructures, and ownership boundaries. Similarly, the Agentic Web highlights intelligent services built on top of existing web platforms, and Web 3.0 mainly advances decentralized data ownership and semantic interoperability. The IoA builds upon but goes beyond these paradigms by integrating autonomy, adaptivity, and economic incentives into a unified, Internet-like fabric for agent collaboration. Table V summarizes key conceptual differences of IoA with MAS, Agentic Web, and Web 3.0.
@@ II-D Emerging IoA Application Scenarios
1) Typical Multi-Agent Frameworks: Recent advancements in LLMs have spurred the development of various agent frameworks capable of autonomous or semi-autonomous task execution through advanced reasoning, planning, and tool integration. AutoGPT [52] implements autonomous goal-driven agents by iteratively refining actions through recursive self-prompting. AutoGen [41] introduces conversational agent networks where multiple LLM-powered agents (e.g., assistants and user proxies) interact dynamically, supporting both human-in-the-loop and fully automated workflows. LangChain [53] provides a modular framework for chaining LLM calls with external tools (e.g., APIs and databases) and memory systems, offering flexibility for customizable agent pipelines. MetaGPT [37] simulates software development teams with role-based agents (e.g., product managers and engineers) that follow structured workflows, enhancing multi-agent collaboration. Other notable frameworks include BabyAGI [54], which focuses on task-driven autonomous iteration, and CAMEL [47], which explores role-playing agent societies with communicative agents. Table VI shows the comparisons of mainstream open-source multi-agent frameworks.
TABLE VI: Comparison of Mainstream Agent Frameworks
@@
MetaGPT[37]
Role-based agents,
structured workflows,
multi-agent collaboration
High efficiency in
complex tasks
Require predefined roles,
less flexible
Automated software
development
LangChain[53]
Modular LLM chaining,
memory, tool integration
Highly customizable,
supports diverse tools
Steeper learning curve,
manual tuning needed
Custom agent pipelines,
RAG systems
AutoGPT[52]
Recursive self-prompting,
goal-driven autonomy
Fully autonomous
task execution
Prone to loops,
high compute cost
General task automation
AutoGen[41]
Conversational MAS,
human-in-the-loop support
Flexible collaboration,
dynamic interactions
Complex setup for
optimal performance
Multi-agent dialogue,
AI teamwork
BabyAGI[54]
Task-driven autonomous
iteration, simple architecture
Lightweight,
easy to deploy
Limited reasoning depth Small-scale automation
CAMEL[47]
Role-playing agent societies,
communicative agents
Simulates human-like
interactions
Less optimized for tool use
Social AI,
research simulations
@@
2) Roadmap: Existing representative industrial practices include computer use, AI smartphones, Anthropic's model context protocol (MCP) [13], and Google's agent-to-agent (A2A) Protocol [9]. (i) Computer-Use Agents (CUAs): As existing Internet services are originally built for humans, CUAs [14] such as OpenAI's Operator mimic human GUI interactions (e.g., mouse clicks and keystrokes) to drive browsers and mobile interfaces, but incur significant screen-scraping overhead that undermines agent efficiency. (ii) AI Smartphones: AI smartphones (e.g., Apple Intelligence) enable cross-App data access and function calls by exposing App APIs to agents, better aligned with agent's strength in processing raw data than GUI emulation. (iii) MCP places large models at the core of the ecosystem, treating the entire Internet as its contextual knowledge base and toolkit. Standardized APIs, dynamic plugin support, and persistent context management allow a single model to orchestrate across diverse services. (iv) A2A protocol aims to form a decentralized agent network where agents autonomously discover, negotiate, and collaborate with each other. Ultimately, IoA should provide (1) seamless interoperability to eliminate data silos, (2) agent-native interfaces for direct efficient inter-agent interactions, and (3) protocol-driven self-organizing cooperation among agents.
3) Real IoA Prototypes and Industrial Case Studies: To illustrate the practical realization of the IoA, we analyze representative academic and industrial IoA implementations.
a) IoA Prototype in [7]. Chen et al. present an IoA prototype to connect heterogeneous virtual agents for collaborative intelligence. It adopts a layered client-server architecture with interaction, data, foundation layers, comprising four core components: an agent registry & discovery module for registration and semantic matching; a task management & orchestration engine for autonomous team formation, subtask allocation, and workflow sequencing; a data/knowledge store that enables context-aware information sharing and local agent memory; and message-protocols that specify agent message formats, group-chat semantics, and a finite-state conversation controller inspired by speech-act theory. By integrating symbolic, LLM-powered, and task-specialized agents, this prototype supports dynamic teaming and conversational flow control rather than fixed, hard-coded pipelines.
b) DAWN [8]. DAWN aims to support globally distributed IoA collaboration via a hierarchical architecture comprising Principal Services Agents and Gateway Services Agents. The Principal Agent plans and orchestrates tasks, while each associated Gateway Agent maintains registries of tools, agents, and applications, exposing them via RESTful endpoints. Upon receiving a task, the Principal Agent delegates subtasks to Gateway Agents, which identify suitable resources and return them for execution. DAWN supports three operational modes: No-LLM mode for highly predictable, deterministic tasks using traditional tools and human-designed workflows; Copilot mode, where agents assist humans with LLM-guided planning and resource orchestration; and LLM agent mode, in which agents operate autonomously, performing reasoning, planning, and execution using LLMs with minimal human intervention.
c) Siemens Industrial Copilot [55]. Unveiled at 2025 in Detroit, Siemens’ Industrial Copilot framework applies autonomous agents across the entire production lifecycle, supporting design optimization, engineering assistance, and predictive maintenance in smart factories. In the design phase, design agents assist engineers with system modeling and workflow optimization. During engineering, planning and engineering agents automate routine programming, system configuration, and cross-team coordination, reducing manual effort and errors. Finally, in maintenance, operation and service agents analyze real-time data to anticipate equipment failures and enable proactive scheduling, supporting predictive maintenance strategies. This multi-agent copilot ecosystem collectively enhances efficiency, reliability, and resilience in industrial automation.
d) Midea Intelligent-Agent Factory [56]. Midea's Jingzhou facility exemplifies an experimental intelligent-agent factory powered by the Factory Brain: a central orchestrator that enables A2A-based multi-agent coordination and leverages an industrial-grade large-model inference engine. The system deploys 14 agents across 38 production scenarios, spanning scheduling, energy management, quality assurance, operations, and maintenance. For instance, in automated quality inspection, quality and R&D agents collaborate via the Factory Brain to implement closed-loop inspections using AI-enabled glasses, reducing first-pass inspection time from 15 minutes to 30 seconds. Embodied humanoid agents such as Meiro are also used to autonomously execute inspections, equipment checks, and safety patrols. Together, these heterogeneous agents demonstrate the potential of coordinated IoA ecosystems to enhance efficiency and flexibility in manufacturing.
@@
4) Applications: The IoA paradigm holds transformative potential across diverse sectors. As shown in Fig. 5, we illustrate its applications in five representative scenarios: smart homes, healthcare, smart grids, smart factories, and smart cities.
@@
  • Scenario 2: Healthcare Robot Coordination within an IoA Subnet in Medical Facilities. In the healthcare context, a dedicated IoA subnet interconnects diverse medical agents, such as surgical robots, anesthesia robots, monitoring robots, and delivery robots, to enable seamless intra-facility coordination. Upon detecting abnormal patient vital signs, a monitoring robot leverages the subnet to discover surgical robots for instrument preparation, anesthesia robots for dosage calculation, and delivery robots for timely medication transport. Each participating agent is assigned a unique digital identity and engages in encrypted, credential-based communication to ensure data integrity and patient confidentiality. By dynamically forming task-specific teams and orchestrating multi-agent workflows, the subnet enhances emergency responsiveness, reduces procedural delays, and improves overall patient safety within medical facilities.